OpenStudy (anonymous):

Simplify the given expression: √-10 times √-8

OpenStudy (anonymous):

sqrt(2*5)*sqrt(2*2*2)=sqrt(2*2*2*2*5)=2*2sqrt(5)=4sqrt(5)

OpenStudy (anonymous):

wait wait is that a negative?

OpenStudy (anonymous):

forget everything I said

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

haha okay

OpenStudy (anonymous):

and is there an i in the answer? like as in imaginary numbers

OpenStudy (anonymous):

ye then it's goes like: i d0Nt KnoW

OpenStudy (anonymous):

\[i \sqrt{10}\times i \sqrt{8}\]

OpenStudy (anonymous):

\[i ^{2}\times \sqrt{80}\]

OpenStudy (anonymous):

\[i ^{2}\times 4\sqrt{5}\]

OpenStudy (anonymous):

\[-4\sqrt{5}\]

OpenStudy (phi):

cblr did it correctly except you for the imaginary part write sqrt(-10)*sqrt(-8) as i * sqrt(10)* i * sqrt(8) or i *i* sqrt(10)* sqrt(8) now you can do cblrtopas idea but you also have a i*i out front (which you know is equal to -1)

OpenStudy (anonymous):

i=sqrt(-1) I think. So, i*sqrt(10)*i*sqrt(8) --> i^2*sqrt(80)=-4*sqrt(2)

OpenStudy (phi):

the idea is break the numbers 10 and 8 into their factors, and pull out pairs of numbers from out of the square root.

OpenStudy (anonymous):

I mean -4*sqrt(5)

OpenStudy (anonymous):

My choices are: A. -4√5 B. 4√5 C. 4i√5 D. -4i√5

OpenStudy (anonymous):

D it up

OpenStudy (anonymous):

Now help me with my problems

OpenStudy (anonymous):

yeah i narrowed it down to A and D. i know it's negative, i just dont know if theres an i or not

OpenStudy (anonymous):

No, it's A seriously

OpenStudy (anonymous):

hahaha okay thanks

OpenStudy (anonymous):

what subject are you working on?

OpenStudy (anonymous):

It's A

OpenStudy (anonymous):

It's my can; y'see....I'm just kidding have a good day sir.